OpenStudy (anonymous):

Write the equation of a line that goes through the points (-4, 5) and (2, -4)

OpenStudy (anonymous):

\[\huge m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] Where m represents your slope. Once you have that use the point - slope form \[\huge y-y _{1}=m(x-x _{1})\] Transform this into \[\huge y = mx+b \]

OpenStudy (anonymous):

Where x1 and y1 are from your given coordinates.

OpenStudy (anonymous):

Thank you for your help.